OpenStudy (anonymous):

HELLLLPPP! How to solve the following using L-hopital rule? 1) limit h-0 = e^h-1/h 2) limit h-0 = ln(1+h)^1/h

OpenStudy (kc_kennylau):

\[\lim_{h\rightarrow0}\frac{e^{h-1}}h\]\[\lim_{h\rightarrow0}\ln(1+h)^{\frac1h}\]

OpenStudy (kc_kennylau):

Are these what you mean?

OpenStudy (anonymous):

yes

OpenStudy (kc_kennylau):

Wait

OpenStudy (kc_kennylau):

Did you mean \(\displaystyle\lim_{h\rightarrow0}\frac{e^h-1}h\) for the first one

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no no i meant the first one

OpenStudy (anonymous):

@kc_kennylau ??

OpenStudy (kc_kennylau):

Then it'd be in the form of 1/0

OpenStudy (anonymous):

i have to prove the first and second question as equal to 1

OpenStudy (kc_kennylau):

Let me ask you again, it's \(\displaystyle\lim_{h\rightarrow0}\frac{e^{h-1}}h\) or \(\displaystyle\lim_{h\rightarrow0}\frac{e^h-1}h\)?

OpenStudy (anonymous):

the first one

OpenStudy (kc_kennylau):

Then it's unsolvable

OpenStudy (kc_kennylau):

You sure it's the first one

OpenStudy (anonymous):

yes i am sure it is the limiting case of natural log

OpenStudy (kc_kennylau):

@mathmale I can't do an impossible question

OpenStudy (anonymous):

L-Hopital's Rule states that if the numerator and denominator both equal zero when the limit is plugged in, then you must take the derivative of the numerator and denominator and then plug in the limit. This will give you the answer. Hope this helps :)

OpenStudy (anonymous):

i know but instead of getting 1 i am getting 0.367 :/

OpenStudy (anonymous):

Can you attach a copy of your work? Maybe I can see what you may be doing wrong

OpenStudy (mathmale):

Forgive me, but I'm going to assume that the problem is actually\[\lim_{h\rightarrow0}\frac{e^{h}-1}h\] If this is the case, then as h approaches zero, e^2 approaches 1 and we end up with 0/0, a perfect case for application of l'Hopital's Rule. Taking the derivative of numerator and denom. separately, we get the lim as h approaches 0 of e^h, which is just 1. @Farheen28: Would you please post a screen shot of the original problem, for verification.

OpenStudy (anonymous):

OMG i am so sorry. it is e^h -1/h

OpenStudy (anonymous):

@kc_kennylau you were right :P

OpenStudy (mathmale):

OK. Problem solved. Note that if the other version were the correct one, l'Hopital's Rule would NOT apply.

OpenStudy (anonymous):

If you click the equation button at the bottom of the text box, you can create the equation in a way where there will be no confusion.

OpenStudy (anonymous):

@Farheen28 are you able to show a copy of your work?

OpenStudy (kc_kennylau):

Well it'd be better to use the "draw" button instead because he/she may not be familiarized with LaTeX.

OpenStudy (anonymous):

\[\ln (1+h)^{1/h}\] this is the other question

OpenStudy (kc_kennylau):

Use the formula \(\Large\log_na^b=b\log_na\).

OpenStudy (anonymous):

@kc_kennylau the question says to use L'Hopital's rule.

OpenStudy (kc_kennylau):

Yes, this is the first step, I'll use the rule later.

OpenStudy (mathmale):

Actually, @wotseit, kc_kenny's approach is absolutely correct.

OpenStudy (anonymous):

i first used that formula but it gave infinity as the answer then what should i do?

OpenStudy (anonymous):

Oh,whoops, I think I misread what you wrote. Sorry!

OpenStudy (kc_kennylau):

No hard feelings

OpenStudy (mathmale):

@Farheen28, what do YOU think? Do as much as you can of this work.

OpenStudy (kc_kennylau):

Show us what you got after you used this formula @Farheen28

OpenStudy (anonymous):

ok wait it'll take some time

OpenStudy (anonymous):

\[\frac{ 1 }{ h }\ln (1+h)\] \[\frac{ 1 }{ h }.\frac{ 1 }{ 1+h }.1\] then i put the limit and got infinity as my answer

OpenStudy (kc_kennylau):

\(\dfrac1h\ln(1+h)\) is in the form of ∞×0...

OpenStudy (kc_kennylau):

You can't apply l'hopital rule to ∞×0

OpenStudy (anonymous):

It seems that instead of taking the derivative of h (which would be the denominator), you left it as 1/h and multiplied that by the derivative of ln(1+h). Don't forget to take the derivative of both the numerator and the denominator when applying L'Hopital's Rule

OpenStudy (anonymous):

are these two different \[\infty \times0\] and \[0\times \infty\] no right? because you can apply l hopital to \[0 \times \infty\]

OpenStudy (kc_kennylau):

And no you can't apply l'hopital to \(0\times\infty\)

OpenStudy (kc_kennylau):

You need to convert \(0\times\infty\) into \(\dfrac00\) before you can use it.

OpenStudy (anonymous):

Where is infinity coming from? Just curious.

OpenStudy (kc_kennylau):

\(\dfrac1h\) itself is infinity

OpenStudy (anonymous):

i am confused

OpenStudy (kc_kennylau):

You can't apply l'hopital to \(\dfrac1h\ln(1+h)\), but you can definitely do so to \(\dfrac{\ln(1+h)}h\).

OpenStudy (kc_kennylau):

Although they are basically the same

OpenStudy (anonymous):

I don't know anything about infinities but if we multiply 1/h by ln (1+h) we get \[\frac{ \ln(1+h) }{ h }\] And then you just apply L'Hopital's rule And kenny beat me again. I'm so slow haha

OpenStudy (anonymous):

could you elaborate how am i going t

OpenStudy (anonymous):

to reach 1 as my answer?

OpenStudy (anonymous):

Apply L'Hopital's rule to \[\frac{ \ln(1+h) }{ h }\] and you will get 1. I can walk you through it if you'd like.

OpenStudy (anonymous):

after differentiating it i will get \[\frac{ 1 }{ 1+h }\] right?

OpenStudy (anonymous):

@wotseit ?

OpenStudy (anonymous):

Yes, you are right. Keep going!

OpenStudy (anonymous):

then when i will put 0 in h i will get 1 as my answer :)

OpenStudy (kc_kennylau):

Exactly :D

OpenStudy (anonymous):

thank you all :D one more thing can i apply l hopital rule when \[0^{0}\], \[0 \times \infty\] and \[1^{\infty}\] ?

OpenStudy (kc_kennylau):

No, l'hopital rule only works for \(\dfrac00\) and \(\dfrac\infty\infty\)

OpenStudy (kc_kennylau):

However, you can convert the above cases into \(\dfrac00\) and then you can use it

OpenStudy (anonymous):

how can i do that?

OpenStudy (kc_kennylau):

\[\Large0^0=e^{\ln0^0}=e^{0\ln0}=e^{\frac{\ln0}{\infty}}=e^{-\frac\infty\infty}\]

OpenStudy (kc_kennylau):

Tell me which step you don't understand if there's any

OpenStudy (anonymous):

3rd

OpenStudy (kc_kennylau):

It's using the formula \(\log a^b=b\log a\)

OpenStudy (kc_kennylau):

If that's what you mean by the "3rd" step

OpenStudy (anonymous):

4th actually

OpenStudy (kc_kennylau):

Because \(\dfrac1\infty=0\)

OpenStudy (kc_kennylau):

\[\Large0^0=e^{\ln0^0}=e^{0\ln0}=e^{\frac1\infty\ln0}=e^{\frac{\ln0}{\infty}}=e^{-\frac\infty\infty}\]

OpenStudy (anonymous):

Think of it like this: infinity is a very large number. If you divide a number by a very large number, you get very close to zero. That's how I like to think of it, at least.

OpenStudy (anonymous):

ah now i get it :) i am sorry i put you in so much trouble :/

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (kc_kennylau):

\[\Large1^\infty=e^{\ln1^\infty}=e^{\infty\ln1}=e^{\infty0}\]

OpenStudy (kc_kennylau):

\[\infty\times0=\frac10\times0=\frac00\]

OpenStudy (anonymous):

thank you :D

OpenStudy (kc_kennylau):

No problem :)

OpenStudy (kc_kennylau):

And as always, \(\infty\times0=\infty\times\dfrac1\infty=\dfrac\infty\infty\)